huff:

Eric had $197 in his savings account before he was paid his weekly salary. His current savings balance is $429. How much money does Eric earn each week? (Form an equation.)

Ferrari:

What do you know about equations like this? I have an answer, but I just want to ask you first.

NoodlesAndRiceYT:

What are you thinking about this question, tell us what you know :D

huff:

well im just really confused on if I need to add or multiply for this equation.

Joe348:

simple 197+ how much =429

Ferrari:

@joe348 wrote:
simple 197+ how much =429
Actually, there is a weekly pay involved, so it is multiplication.

huff:

so would it be 197+n=429?

Joe348:

@huff wrote:
so would it be 197+n=429?
yes

NoodlesAndRiceYT:

No multiplication should be involved.

Joe348:

Bruh no there's not

Ferrari:

Well, there is a weekly price. Also, notice that it says "each" which signals multiplication...

Joe348:

His total is 429 and he has 197 he needs to add more money to get to that price

huff:

so which one is it

Joe348:

huff you can listen to them or me your choice bro but ik im right

huff:

197n=429 or 197+n=429

huff:

okay thank you guys for the help!

Ferrari:

However, I will just go with what they say... IG. P = $429 - $197 or... $197 + P = $429.

ILOVESPAGHETTI:

I think it's multiplication.

huff:

adding was right by the way

NoodlesAndRiceYT:

there will be no Multiplication because its not talking about more then one week, just talking about that one week itself

ILOVESPAGHETTI:

@noodlesandriceyt wrote:
there will be no Multiplication because its not talking about more then one week, just talking about that one week itself
ah, makes sense. this is why i stay in english

NoodlesAndRiceYT:

@noodlesandriceyt wrote:
there will be no Multiplication because its not talking about more then one week, just talking about that one week itself
Make sense @ferrari ?

Ferrari:

Yes. I have some issues in my mental health where I tend to overthink things... Thank you!